# Newton numbers, vanishing polytopes and algebraic degrees

### v1

@@ 6.1 Preliminaries: Dual defective sets
In this subsection, we recall the Furukawa-Ito classification of dual defective sets from [FI21].
The set $\mathbf{A}\subset\mathbb{Z}^{n}$ is called dual defective if the projective dual set to the corresponding projectively toric variety $X_{\mathbf{A}}\subset\mathbb{C}P^{|\mathbf{A}|}$ is not a hypersurface (see e.g. [FI21]). The projective dual set to $X_{\mathbf{A}}$ is known as the Gelfand-Kapranov-Zelevinsky $\mathbf{A}$-discriminant surface (see [GKZ94, Chapter 9, Proposition 1.1.]). The set $\mathbf{A}$ is called dual defective if and only if the $\mathbf{A}$-discriminant is equal to $1$. Denote by $\mathbb{C}[\mathbf{A}]$ the set of Laurent polynomials with exponents of their monomials contained in $\mathbf{A}$.
@@
Let us recall the classification of dual defective sets in terms of Cayley sums from [FI21]. For a different description of the dual defective sets using so called ‘‘circuits’’ see [Es17, Corollary 3.20]. For $\mathbf{A}\subset\mathbb{Z}^{n}$ we denote by $\langle\mathbf{A}-\mathbf{A}\rangle$ the subgroup of $\mathbb{Z}^{n}$ generated by the set $\{a_{1}-a_{2}:a_{1},a_{2}\in\mathbf{A}\}$ and say that $\mathbf{A}\subset\mathbb{Z}^{n}$ is spanning, if $\langle\mathbf{A}-\mathbf{A}\rangle=\mathbb{Z}^{n}$.
@@ 6.2 Preliminaries: Thin polytopes
Thin simplices were defined in [GKZ94, 11.4.B], where they were classified up to dimension $2$. In [GKZ94] they define thin simplices through the Newton number (here it is the alternating sum of volumes of all faces of the polytope including the empty face). Note that this Newton number has nothing in common with the local $h^{*}$-polynomial an for arbitrary polytope, for instance, it can be negative (it equals $-1$ for the cube $[0,1]^{3}$), hence such definition cannot be extended to arbitrary polytopes.
@@
Recall that the degree $\deg(P)$ of the polytope $P$ is the degree of its $h^{*}$-polynomial, the codegree codeg$(P)$ is the smallest integer $\lambda\geq 1$ such that the $\lambda$-th dilate $\lambda P$ of $P$ contains a lattice point in its relative interior. By convention, a point has codegree $1$. It follows from Ehrhart-MacDonald reciprocity [Mcd71] that
@@
[BKN24, Theorem 4.3.] Let $P$ be a $3$-dimensional lattice polytope. Then $P$ is thin if and only if up to isomorphism
@@
Polytope $P$ is called trivially thin if $\dim(P)\geq 2\deg(P)$
@@
(see e.g. [BKN24, Proposition 3.14.]) Let $P\subset\mathbb{R}^{n}$ and $Q\subset\mathbb{R}^{m}$ be lattice polytopes. Then
@@
Question 6.16.
(already answered) [BKN24, Question 3.16.]
@@
Note that in dimension $3$ every thin polytope is either a lattice pyramid or trivially thin. But the situation is not the same for arbitrary dimension. In [Kur24b, Corollary 2] (using so-called ‘‘Lawrence twist’’) there is obtained an example of spanning thin polytope in dimension $5$, which is not a join. We recall it in Example 6.28 and show that it is a $B_{2}$-polytope. An up-to-date question concerning the classification is as follows.
Question 6.17.
@@
Recall that the width of a lattice polytope is the minimal lattice distance (number of parallel lattice hyperplanes inside plus $1$) between parallel hyperplanes, so that the polytope is contained between them. Let us also mention another up-to-date question concerning thin polytopes (though our technique provides no insight concerning it since the width is equal to one if and only if the polytope is a non-trivial Cayley sum).
Question 6.18.
([BKN24, Question 3.19] and [Kur24b, Conjecture 6,7]) What is the possible width of thin polytopes?
@@
Question 6.20.
Can we classify thin polyhedral subdivision of polytopes (or cones)?
The question is not answered even for triangulations of simplices. The answer is known for $2$- and $3$-dimensional simplices (see [MGPSS20]). For more results concerning thin triangulations of simplices (also called triangulations with vanishing local $h$-polynomials), see [LPS22] and [LPS23]. The classification problem is significantly motivated by the monodromy conjecture (see e.g. [LPS22] and [Ve]). It is also important for the Arnold’s monotonicity problem [Sel24] (the classification of thin triangulation in the cones $\mathbb{R}^{3}_{\geq 0}\oplus\mathbb{R}^{1}$ and $\mathbb{R}^{2}_{\geq 0}\oplus\mathbb{R}^{2}$ conjectured in [Sel24, Conjecture 2.28] leads to the complete solution of the Arnold’s monotonicity problem in dimension $5$, see [Sel24, Theorem 3.6] and the remark below it). We believe that the same motivation can be redirected to the problem of classification of thin polytopes.
@@
We believe that polyhedral subdivisions of the $B_{k}$-polytopes provide non-trivial examples of thin polyhedral subdivisions.
6.3 Properties of the $B_{k}$-polytopes
By the $B_{k}$-Theorem 1.3 any $B_{k}$-polytope in $\mathbb{R}^{n}$ based on $\mathbb{R}^{k}_{\geq 0}$ is negligible in the cone $\mathbb{R}^{k}_{\geq 0}\oplus\mathbb{R}^{n-k}$. By Theorem 4.6 the negligible polytopes in $\mathbb{R}^{k}_{\geq 0}\oplus\mathbb{R}^{n-k}$ are thin. Thus $B_{k}$-polytopes are thin. Let us establish some properties of $B_{k}$-polytopes showing that they are non-trivial examples of thin polytopes.
@@ 6.3 Properties of the $B_{k}$-polytopes
Using the language of [Nil20] the second part of the proposition claims that $\text{codeg}(P_{0}\ast P_{1}\dots\ast P_{k})=k+1$ if and only if we have the following inequality for the mixed codegree $\text{mcd}(P_{0},\dots,P_{k})\leq k+1$, i.e. $\text{mcd}(P_{0},\dots,P_{k})\neq k+2$.
@@
Note that the projection of $\lambda P$ on $\mathbb{R}^{k}$ contains no interior lattice points if $\lambda\leq k$ and this proves the first part of the proposition. The projection of $(k+1)P$ on $\mathbb{R}^{k}$ contains a unique interior lattice point $(1,\dots,1)$. The fiber over it is exactly the polytope $P_{0}+\dots+P_{k}$ and this proves the second part of the proposition. ∎
@@
Assume that $P$ is a join of $Q_{1}$ and $Q_{2}$. Since $P_{0}$ is not a join, then it is contained either in $Q_{1}$ or in $Q_{2}$. Without loss of generality, assume that $P_{0}\subset Q_{1}$. Then, since $\text{aff}(P_{0})=\mathbb{R}^{n-k}$, each $P_{i},i\geq 1$ is either contained in $Q_{1}$ or contained in $Q_{2}$. Some of them must be contained in $Q_{2}$ since $Q_{2}\neq\emptyset$. Recall that $\langle P_{i}-P_{i}\rangle\subset\text{aff}(P)$, so it is not a join. ∎
@@
From Corollary 1.7 we obtain that this polytope is thin. By Corollaries 6.23 and 6.26 this polytope is neither trivially thin nor a join.
@@
Their sum contains the interior lattice point $(0,1,0)$, thus, by Corollary 6.23 the polytope $P$ is not trivially thin. By Corollary 6.26 it is not a join.
6.4 Proof of the $B_{k}$-Theorem
In this section we prove the $B_{k}$-Theorem 1.3.
We split the proof into two parts. First, we check that $B_{k}$-polytopes are indeed negligible and this is the easy part. Second, using the Furukawa-Ito classification of dual defective sets, we prove that if a polytope $P$ is negligible, then it is a $B_{k}$-polytope.
6.4.1 $B_{k}$-polytopes are negligible
@@
(Simple part of the $B_{k}$-theorem) Consider the cone $C=\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$. Then any $B_{k}$-polytope $P\subset\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ based on a coordinate subspace $\mathbb{R}^{k}_{\geq 0}\subset\mathbb{R}^{m}_{\geq 0}$ is negligible.
@@
First let us prove the lemma for $B_{k}$-polytopes in $C=\mathbb{R}^{k}_{\geq 0}\oplus\mathbb{R}^{n-k}$. Applying Lemma 5.10 to $P_{0}\ast P_{1}\ast\cdots\ast P_{k}$ and $P_{1}\ast\cdots\ast P_{k}$ we obtain:
$\nu_{Q}(P_{0}\ast P_{1}\ast\cdots\ast P_{k})=\sum\limits_{\begin{subarray}{c}a_{0}\geq 0,a_{1}>0,\dots,a_{k}>0\\ a_{0}+\dots+a_{k}=n-k\end{subarray}}MV(\underbrace{P_{0},\dots,P_{0}}_{a_{0}},\dots,\underbrace{P_{k},\dots,P_{k}}_{a_{k}})$
Recall that by the definition of $B_{k}$-polytopes, we have $\dim(P_{1}+\dots+P_{k})<k$, so by Proposition 5.5 each summand of the above sum is equal to 00. Thus, we obtained that the Newton number vanishes:
@@
Now we pass to the arbitrary case $C=\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m},m\geq k$. Let us decompose the cone $C$ in the following way $C=\mathbb{R}^{k}_{\geq 0}\oplus\mathbb{R}^{m-k}_{\geq 0}\oplus\mathbb{R}^{n-m}$. To avoid confusion, denote the first summand by $L_{k}=\mathbb{R}^{k}_{\geq 0}$ and the last by $L_{n-m}=\mathbb{R}^{n-m}$. For $J\subset\{k+1,\dots,m\}$, denote by $E_{J}\subset\mathbb{R}^{n}$ the vector space generated by the corresponding coordinate vectors. Denote by $C_{J}\subset\mathbb{R}^{n},\dim C_{J}=n+k-m+|J|$ the cone $L_{k}+E_{J}+L_{n-m}$. Then we have:
@@
Note that each summand is a $B_{k}$-polytope in a cone $\mathbb{R}^{k}_{\geq 0}\oplus\mathbb{R}^{n-m+|J|}$ based on $\mathbb{R}^{k}_{\geq 0}$. Hence all these summands are equal to 00 as we checked in the beginning of the proof of the lemma.
6.4.2 Negligible polytopes are $B_{k}$-polytopes
Consider a negligible polytope $P\subset C=\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$. Denote by $\mathbf{P}\subset\mathbb{Z}^{m}_{\geq 0}\oplus\mathbb{Z}^{n-m}$ a lattice set such that $\operatorname{Conv}(\mathbf{P})=P$ (e.g. the set of the vertices of $P$ or $P\cap\mathbb{Z}^{n}$).

### v2

@@ 6.1 Preliminaries: Dual defective sets
The set $\mathbf{A}\subset\mathbb{Z}^{n}$ is called dual defective if the projective dual set to the corresponding projectively toric variety $X_{\mathbf{A}}\subset\mathbb{C}P^{|\mathbf{A}|}$ is not a hypersurface (see e.g. [FI21]). The projective dual set to $X_{\mathbf{A}}$ is known as the Gelfand-Kapranov-Zelevinsky $\mathbf{A}$-discriminant surface (see [GKZ94, Proposition 1.1 from Chapter 9]). Equivalently, the set $\mathbf{A}$ is dual defective if and only if the $\mathbf{A}$-discriminant is equal to $1$.
In this paper we need the Furukawa-Ito classification of dual defective sets from [FI21], so we recall it in this subsection. First, let us briefly recall all the known combinatorial descriptions of dual defective sets (see [CD22] for the overview).
There are four descriptions combinatorial of dual defective sets in terms of:
  1. 1.
    ranks of certain matrices associated to the vectors in the kernel of $\mathbf{A}$ [DFS07].
  2. 2.
    a matroid associated to the Gale transform of $\mathbf{A}$ [CC07].
  3. 3.
    existence of (iterated) circuits in $\mathbf{A}$ [Es18].
  4. 4.
    existence of certain Cayley decomposition of $\mathbf{A}$ called Furukawa-Ito decomposition [FI21] (see Theorem 6.5 in our paper).
These combinatorial descriptions are equivalent since they all describe the dual defective sets. In [CD22] all these approaches are compared from the combinatorial point of view. It is shown that the items 1-3 are equivalent using linear-algebraic methods which require only the combinatorial definitions. However for the Furukawa-Ito classification (combinatorially) it is only proved that if a set has Furukawa-Ito decomposition then it also satisfies items 1-3, but not vise versa (see [CD22, Remark 3]).
Question 6.1.
Can we prove combinatorially (using linear-algebraic methods) that descriptions from [DFS07], [CC07], [Es18] (items 1-3) of dual defective sets imply existence of Furukawa-Ito decomposition (item 4)?
The latter question appears to be surprisingly non-trivial. For instance, our proof of $B_{k}$-theorem 1.3 significantly relies on the Furukawa-Ito classification of dual defective sets and we do not know a proof using other combinatorial descriptions of dual defective sets.
Recall that $\mathbb{C}[\mathbf{A}]$ is the vector space of Laurent polynomials with supports contained in the finite lattice set $\mathbf{A}$.
@@
Let us recall the classification of dual defective sets in terms of Cayley sums from [FI21]. Recall that for $\mathbf{A}\subset\mathbb{Z}^{n}$ we denote by $\langle\mathbf{A}-\mathbf{A}\rangle$ the subgroup of $\mathbb{Z}^{n}$ generated by the set $\{a_{1}-a_{2}:a_{1},a_{2}\in\mathbf{A}\}$ and say that $\mathbf{A}\subset\mathbb{Z}^{n}$ is spanning, if $\langle\mathbf{A}-\mathbf{A}\rangle=\mathbb{Z}^{n}$.
@@ 6.2 Preliminaries: Thin polytopes
Thin simplices were defined in [GKZ94, Chapter 11.4.B], where they were classified up to dimension $2$. In [GKZ94] they define thin simplices through the Newton number (here it is the alternating sum of volumes of all faces of the polytope including the empty face). Note that this Newton number has nothing in common with the local $h^{*}$-polynomial for arbitrary polytopes, for instance, it can be negative (it equals 1-1 for the cube $0,1]^{3}$), hence such definition cannot be extended to arbitrary polytopes.
@@
Recall that the degree $\deg(P)$ of a polytope $P$ is the degree of its $h^{*}$-polynomial. The codegree $\text{codeg}(P)$ is the smallest integer $\lambda\geq 1$ such that the $\lambda$-th dilate $\lambda P$ contains a lattice point in its relative interior. By convention, a point has codegree $1$. It follows from Ehrhart-MacDonald reciprocity [Mcd71] that
@@
[BKN24, Theorem 4.3] Let $P$ be a $3$-dimensional lattice polytope. Then $P$ is thin if and only if up to isomorphism
@@
Remark 6.12.
For $k\geq 2$ every $B_{k}$-polytope in $\mathbb{R}^{k+1}$ is either a lattice pyramid or a Lawrence prism. Moreover, every Lawrence prism in dimension $k+1\geq 3$ is isomorphic to such a $B_{k}$-polytope. In dimension $3$ all thin polytopes are equivalent to $B_{k}$-polytopes (for $k=1$ or $2$). In dimension $4$ there are already other examples of thin polytopes (including simplices which are not Cayley sums at all, see [Kur24a]). So $B_{k}$-polytopes can be considered as a generalization of lattice pyramids and Lawrence prisms.
@@
A polytope $P$ is called trivially thin if $\dim(P)\geq 2\deg(P)$.
@@
(see e.g. [BKN24, Proposition 3.14]) Let $P\subset\mathbb{R}^{n}$ and $Q\subset\mathbb{R}^{m}$ be lattice polytopes. Then
@@
Question 6.18.
(already answered in [Kur24b, Corollary 2]) [BKN24, Question 3.16]
@@
Note that in dimension $3$ every thin polytope is either a lattice pyramid or trivially thin. In [Kur24b, Corollary 2] (using so-called Lawrence twist) there is obtained an example of spanning thin polytope in dimension $5$, which is not a join. We recall it in Example 6.30 and show that it is a $B_{2}$-polytope. An up-to-date question concerning the classification is as follows.
Question 6.19.
@@
Another up-to-date question is as follows.
Question 6.20.
([BKN24, Question 3.19] and [Kur24b, Conjectures 6, 7]) What is the possible width of thin polytopes?
@@
Question 6.22.
Can we classify thin polyhedral subdivisions of polytopes (or cones)?
The question is not answered even for triangulations of simplices. The answer is known for $2$- and $3$-dimensional simplices (see [dMGPSS20]). For more results concerning thin triangulations of simplices (also called triangulations with vanishing local $h$-polynomials), see [LPS23a] and [LPS23b].
The classification problem is significantly motivated by the monodromy conjecture (see e.g. [LPS23a] and [Ve]). It is also important for Arnold’s monotonicity problem. Namely, the classification of thin triangulations of the cones $\mathbb{R}^{3}_{\geq 0}\oplus\mathbb{R}^{1}$ and $\mathbb{R}^{2}_{\geq 0}\oplus\mathbb{R}^{2}$, conjectured in [Sel24, Conjecture 2.28], would lead to the complete solution of Arnold’s monotonicity problem in dimension $5$ (see [Sel24, Theorem 3.6] and the remark below it). We believe that the same motivation can be redirected to the problem of classification of thin polytopes.
@@
We believe that polyhedral subdivisions of $B_{k}$-polytopes provide non-trivial examples of thin polyhedral subdivisions.
6.3 Properties of $\mathbf{B_{k}}$-polytopes
By the $B_{k}$-Theorem 1.3 any $B_{k}$-polytope in $\mathbb{R}^{n}$ based on $\mathbb{R}^{k}_{\geq 0}$ is negligible in the cone $\mathbb{R}^{k}_{\geq 0}\oplus\mathbb{R}^{n-k}$. By Theorem 4.6 the negligible polytopes in $\mathbb{R}^{k}_{\geq 0}\oplus\mathbb{R}^{n-k}$ are thin. Thus $B_{k}$-polytopes are thin. Let us establish some properties of $B_{k}$-polytopes showing that they are non-trivial examples of thin polytopes. We compare $B_{k}$- and $B$-polytopes from [ELT22, Definition 1.4] in §6.3.1.
@@ 6.3 Properties of $\mathbf{B_{k}}$-polytopes
Using the language of [Nil20], the second part of the proposition claims that $\text{codeg}(P_{0}\ast P_{1}\dots\ast P_{k})=k+1$ if and only if we have the following inequality for the mixed codegree $\text{mcd}(P_{0},\dots,P_{k})\leq k+1$, i.e. $\text{mcd}(P_{0},\dots,P_{k})\neq k+2$.
@@
Note that the projection of $\lambda P$ onto $\mathbb{R}^{k}$ contains no interior lattice points if $\lambda\leq k$, which proves the first part of the proposition. The projection of $(k+1)P$ onto $\mathbb{R}^{k}$ contains a unique interior lattice point $(1,\dots,1)$. The fiber over it is exactly the polytope $P_{0}+\dots+P_{k}$ and this proves the second part of the proposition. ∎
@@
Assume that $P$ is a join of $Q_{1}$ and $Q_{2}$. Since $P_{0}$ is not a join, then it is contained either in $Q_{1}$ or in $Q_{2}$. Without loss of generality, assume that $P_{0}\subset Q_{1}$. Then, since $\text{aff}(P_{0})=\mathbb{R}^{n-k}$, each $P_{i}$, $i\geq 1$, is either contained in $Q_{1}$ or contained in $Q_{2}$. Some of them must be contained in $Q_{2}$ since $Q_{2}\neq\emptyset$. Recall that $\langle P_{i}-P_{i}\rangle\subset\text{aff}(P)$, so it is not a join. ∎
@@
From Corollary 1.7 we obtain that this polytope is thin. By Corollaries 6.25 and 6.28 this polytope is neither trivially thin nor a join.
@@
Their sum contains the interior lattice point $(0,1,0)$; thus, by Corollary 6.25, the polytope $P$ is not trivially thin. By Corollary 6.28 it is not a join.
6.3.1 Comparison of $\mathbf{B_{k}}$-polytopes and $\mathbf{B}$-polytopes
Let us compare $B_{k}$- and $B$-polytopes from [ELT22, Definition 1.4]. Any $B_{k}$-polytope is a $B$-polytope, but not vice versa. First let us recall the definition of $B$-polytopes.
Definition 6.31.
  1. 1.
    A lattice simplex in $\mathbb{R}_{\geq 0}^{m}\oplus\mathbb{R}^{n-m}$ with the standard coordinate system $v_{1},\dots,v_{m},w_{m+1},\dots,w_{n}$ is called a $\mathbf{B_{1}}$-simplex with respect to the $i$-th coordinate if one of its vertices lies in the plane $v_{i}=1$ and all other vertices lie in the plane $v_{i}=0$.
  2. 2.
    A lattice polytope in $\mathbb{R}_{\geq 0}^{m}\oplus\mathbb{R}^{n-m}$ is called a $\mathbf{B}$-polytope, if every lattice simplex that it contains is a $B_{1}$-simplex.
The following remarks admit straightforward verification.
Remark 6.32.
Every $B_{k}$-polytope in $\mathbb{R}_{\geq 0}^{m}\oplus\mathbb{R}^{n-m}$ is a $B$-polytope.
Remark 6.33.
A convenient polytope $P\subset\mathbb{R}_{\geq 0}^{m}\oplus\mathbb{R}^{n-m}$ with $n\leq 3$ is a $B$-polytope if and only if it is a $B_{k}$-polytope for some $k\leq m$.
Starting from dimension $4$ the $B_{k}$- and $B$-polytopes are different.
Example 6.34.
The convex hull of the following set of points in dimension $4$ is a $B$-polytope but not a $B_{k}$-polytope (in $\mathbb{R}^{4}_{\geq 0}$).
$O$ AA $B$ $C$ $F$ $G$ $H$
0 $e_{1}$ $e_{2}$ $e_{1}+e_{2}$ $e_{3}$ $e_{4}$ $e_{3}+e_{4}$
Figure 7: Example of a $B$-polytope which is not a $B_{k}$-polytope
6.4 Proof of the $\mathbf{B_{k}}$-Theorem
In this section we prove $B_{k}$-Theorem 1.3.
We split the proof into two parts. First, we check that $B_{k}$-polytopes are indeed negligible and this is the easy part. Second, using the Furukawa-Ito classification of dual defective sets, we prove that every negligible polytope is a $B_{k}$-polytope.
6.4.1 $\mathbf{B_{k}}$-polytopes are negligible
@@
(Simple part of $B_{k}$-theorem) Consider the cone $C=\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$. Then any $B_{k}$-polytope $P\subset\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ based on a coordinate subspace $\mathbb{R}^{k}_{\geq 0}\subset\mathbb{R}^{m}_{\geq 0}$ is negligible ($\nu_{C}(P)=0$).
@@
First, let us prove the lemma for $B_{k}$-polytopes in $C=\mathbb{R}^{k}_{\geq 0}\oplus\mathbb{R}^{n-k}$. Applying Lemma 5.11 to $P_{0}\ast P_{1}\ast\cdots\ast P_{k}$ and $P_{1}\ast\cdots\ast P_{k}$ we obtain:
$\nu_{Q}(P_{0}\ast P_{1}\ast\cdots\ast P_{k})=\sum\limits_{\begin{subarray}{c}a_{0}\geq 0,a_{1}>0,\dots,a_{k}>0\\ a_{0}+\ldots+a_{k}=n-k\end{subarray}}\operatorname{MV}(\underbrace{P_{0},\dots,P_{0}}_{a_{0}},\dots,\underbrace{P_{k},\dots,P_{k}}_{a_{k}})$
Recall that by the definition of $B_{k}$-polytopes, we have $\dim(P_{1}+\dots+P_{k})<k$; by Proposition 5.6 each summand in the above sum equals 0. Thus, the Newton number vanishes:
@@
Now, we consider the general case $C=\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$ with $m\geq k$. Let us decompose the cone $C$ as follows: $C=\mathbb{R}^{k}_{\geq 0}\oplus\mathbb{R}^{m-k}_{\geq 0}\oplus\mathbb{R}^{n-m}$. To avoid confusion, denote the first summand by $L_{k}=\mathbb{R}^{k}_{\geq 0}$ and the last summand by $L_{n-m}=\mathbb{R}^{n-m}$. For $J\subset\{k+1,\dots,m\}$, denote by $E_{J}\subset\mathbb{R}^{n}$ the vector space generated by the corresponding coordinate vectors. Denote by $C_{J}\subset\mathbb{R}^{n}$ the cone $L_{k}+E_{J}+L_{n-m}$, which has dimension $\dim C_{J}=n+k-m+|J|$. Then we have:
@@
Note that each summand is a $B_{k}$-polytope in the cone $\mathbb{R}^{k}_{\geq 0}\oplus\mathbb{R}^{n-m+|J|}$ based on $\mathbb{R}^{k}_{\geq 0}$. Hence, all these summands are equal to 0, as established at the beginning of the proof.
6.4.2 Negligible polytopes are $\mathbf{B_{k}}$-polytopes
Consider a negligible polytope $P\subset C=\mathbb{R}^{m}_{\geq 0}\oplus\mathbb{R}^{n-m}$. Denote by $\mathbf{P}\subset\mathbb{Z}^{m}_{\geq 0}\oplus\mathbb{Z}^{n-m}$ a lattice set such that $\operatorname{Conv}(\mathbf{P})=P$ (e.g., the set of vertices of $P$ or $P\cap\mathbb{Z}^{n}$).